8=X/3-8
I KNOW U ADD 8 FROM BOTH SIDES BUT THEN WHAT

Answer by checkley77(12844)

8=X/3-8 ADD 8
16=X/3 CROSS MULTIPLY
X=16*3
X=48 ANS.
PROOF:
8=48/3-8
8=16-8
8=8